OpenStudy (destinymasha):

Solve. -s/3 is greater than or equal to 6 A. s ≥ –2 B. s ≤ –2 C. s ≥ –18 D. s ≤ –18

OpenStudy (anonymous):

\[\frac{ -s }{ 3 } \ge 6\]\[-s \ge 18\]\[s \le -18\]

OpenStudy (destinymasha):

ok, so you multiply 6 x 3 to get 18. Thanks

OpenStudy (anonymous):

Yes, because you need to isolate the unknown :)

OpenStudy (destinymasha):

[: That s true. Now i remember math class